OpenStudy (anonymous):

Solving Rational Equations: 1/v^2= 1- 2v^2 + 2v -144/v^2 Please help! Confused, please show work

OpenStudy (shamil98):

Get the constants on one side and the terms with variables in them on the other, then combine the like terms. Could you attempt this?

OpenStudy (anonymous):

So, multiple by v^2

OpenStudy (anonymous):

but then I'm faced with v^2 times 2v... how does that work?

OpenStudy (anonymous):

does it reduce to 1= v^2 -2v- 142?

OpenStudy (shamil98):

1/v^2= 1- 2v^2 + 2v -144/v^2 1/v^2 + 2v^2 - 2v + 144/v^2 = 1 145/v^2 + 2v^2 - 2v = 1 145 + 2v^4 - 2v^3 = 1 2v^4 - 2v^3 + 144 = 0

OpenStudy (anonymous):

Do I solve as a

OpenStudy (anonymous):

quadratic?

OpenStudy (anonymous):

Cause I need to find "V"

OpenStudy (shamil98):

v^4 - v^3 + 72 = 0 Now, factor it out. and then solve for v. You will get complex solutions.

OpenStudy (anonymous):

I don't think that right how Im doing it, cause my class does this. I'll attach my work for another problem

OpenStudy (anonymous):

\[\frac{ 6 }{ b ^{2} } + \frac{ 6b-48 }{ 7b ^{2} }=\frac{ 1 }{ 7b ^{2} }\]

OpenStudy (anonymous):

So then I multiplied both sides by 7b^2

OpenStudy (anonymous):

and it reduced to "6/7+6b-48=1"

OpenStudy (anonymous):

then I solved algebraically to get 8 and 3/6

OpenStudy (anonymous):

@shamil98 help please

OpenStudy (shamil98):

6/b^2 + 6b -48/7b^2 = 1/7b^2 42/b^2 + 6b - 48/7b^2 = 1/7b^2 6(b-1)/7b^2 = 1/7b^2

OpenStudy (shamil98):

42/7b^2*